OpenStudy (anonymous):

Determine the following for the quadratic function f(x) = 2 x^2 + 5 x - 7: Find the coordinates of the vertex (h,k)

OpenStudy (anonymous):

I already know h, by the way, just not k.

OpenStudy (anonymous):

k = f(h) so u can calculate k

OpenStudy (anonymous):

I know. I must have made an arithmetical error. I just can't seem to notice what it is.

OpenStudy (anonymous):

here's what I have so far: f(5/4) = 2 (5/4)^2 + 5 (5/4) -7

OpenStudy (anonymous):

25/8+25/4-7 = 75/8 -7 = 19/8 (which is apparently wrong)

OpenStudy (anonymous):

Can you see any errors within my calculations?

OpenStudy (anonymous):

Hi there, qpHalcy0n

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

I'm not a calculator, give me a sec ;)

OpenStudy (anonymous):

Naturally ;)

OpenStudy (anonymous):

i think your calculations are correct

OpenStudy (anonymous):

Well, first off, your signs are incorrect. h = -b/2a

OpenStudy (anonymous):

Um...but the website we use approved of my definition of h.

OpenStudy (anonymous):

And...I mean, after all a does equal 2 and b 5, so...why wouldn't -5/4 be correct?

OpenStudy (anonymous):

They are incorrect or you wrote the problem down incorrectly. I'm going with the latter.

OpenStudy (anonymous):

Oh, wait...nvm

OpenStudy (anonymous):

So use -5/4 and see what you get now.

OpenStudy (anonymous):

-81/8, which is correct. Thank you for making me aware of my own stupid mistakes!!

OpenStudy (anonymous):

No prob.

OpenStudy (anonymous):

Have a good day/night!!